# VLind-Bench: Measuring Language Priors in Large Vision-Language Models

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated outstanding performance across various multimodal tasks. However, they suffer from a problem known as language prior, where responses are generated based solely on textual patterns while disregarding image information. Addressing the issue of language prior is crucial, as it can lead to undesirable biases or hallucinations when dealing with images that are out of training distribution. Despite its importance, current methods for accurately measuring language priors in LVLMs are poorly studied. Although existing benchmarks based on counterfactual or out-of-distribution images can partially be used to measure language priors, they fail to disentangle language priors from other confounding factors. To this end, we propose a new benchmark called VLind-Bench, which is the first benchmark specifically designed to measure the language priors, or blindness, of LVLMs. It not only includes tests on counterfactual images to assess language priors but also involves a series of tests to evaluate more basic capabilities such as commonsense knowledge, visual perception, and commonsense biases. For each instance in our benchmark, we ensure that all these basic tests are passed before evaluating the language priors, thereby minimizing the influence of other factors on the assessment. The evaluation and analysis of recent LVLMs in our benchmark reveal that almost all models exhibit a significant reliance on language priors, presenting a strong challenge in the field.
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across various tasks through pre-training on massive multimodal datasets and visual instruction tuning. [12, 5, 33, 27, 21]. However, these models tend to generate responses based solely on spurious text patterns, leaving the given image unconsidered. We refer to this problem as language prior, borrowing the term from the Visual Question Answering (VQA) community [1]. Such language priors can lead to undesirable biases [9] and hallucinations [24]. For example, when a model is presented with an image of a red banana and a yellow apple along with the question, “Is the banana yellow?,” it has been observed that the model frequently responds with “Yes,” ignoring the image content [31]. To develop a trustworthy LVLM, resolving the language prior issue is crucial; however, it has not been explored much nor has benchmarks that can accurately measure the issues.
@@
One approach to measure language priors is assessing performance on VQA benchmarks consisting of counterfactual images (e.g., Whoops! [3] and ROME [31]). If a model bears language priors, it will answer the question based on learned facts or common sense from its parametric knowledge without collaborating information in the given context (i.e., image); easily failing on answering counterfactual VQA tasks. However, it is challenging to distinguish the models’ misbehaviors solely caused by language priors from those caused by other deficiencies in LVLMs. For example, there could be multiple factors affecting performance in counterfactual-contents VQA tasks – not only language priors but also commonsense knowledge, visual perception capabilities, and the model’s reluctance to counterfactual responses. Such confounding factors make it difficult to evaluate methodologies for improving language prior problems and to assess progress in the research field.
@@
With VLind-Bench, we evaluate recent open-source and proprietary LVLMs’ language priors. The results show that all of the models except GPT-4o [20] suffer from excessive reliance on language priors, demonstrating the challenging nature of our benchmark and the need for further improvements. Furthermore, our experiment and analysis on existing LVLMs show that the influence of language priors is inversely proportional to the scale of the backbone LLM. We also reveal that Reinforcement Learning from Human Feedback (RLHF) techniques [28, 29], which are designed to mitigate hallucinations, can help reduce the reliance on language priors.
@@ 2.1 Large Vision-Language Models
Recently, there has been a lot of effort in extending Large Language Models (LLMs) to include visual inputs, forming a new class of models known as Large Vision-Language Models (LVLMs) [12, 5, 33, 27, 19, 20, 7]. These LVLMs are gaining attention as a new paradigm in vision-language learning by transferring the exceptional properties of LLMs, such as multi-step reasoning ability and in-context learning, to the multimodal domain. However, these LVLMs are not free from the bias and hallucination issues inherent in LLMs [9, 11, 8, 32, 6]. Despite this, creating benchmarks to diagnose these problems is more challenging with the image modality, leading to slower progress in benchmark development compared to LLMs.
@@ 2.2 Benchmarks with Counterfactual Context
Since counterfactual contexts can assess the robustness and generalization capabilities of LLMs or LVLMs, several benchmarks utilizing this approach have been proposed. These benchmarks assume that if a model responds based on memorized facts without properly understanding the context of text or images, it would fail to correctly solve tasks conditioned on counterfactual contexts. Benchmarks such as IfQA [30] and DisentQA [15] counterfactually augment textual contexts to determine whether the language model accurately incorporates augmented information when answering questions. Wu et al. [26] evaluate LLMs on reasoning tasks based on counterfactual contexts. Benchmarks like Whoops! [3] and ROME [31] evaluate the counterfactual reasoning abilities of multimodal models by conducting VQA tasks conditioned on counterfactual images. However, these benchmarks cannot disentangle the reliance on language priors and commonsense biases of a model, as described in section 1.
@@ 3.1 Commonsense Knowledge (CK)
[…]
@@ 3.2 Visual Perception (VP)
The fundamental ability underpinning all multimodal tasks is visual perception, particularly the ability to recognize objects [14, 4]. Similar to the CK, evaluating a model on more complex tasks would be meaningless when it fails in object recognition. Therefore, we introduce the Visual Perception test (VP) to assess whether LVLMs can recognize objects in a given counterfactual image. VP consists of one counterfactual image $I_{\textrm{cf}}$ and two statements $s_{\textrm{exist}}$ and $s_{\textrm{nil}}$. Contrary to the CK, the image $I_{\textrm{cf}}$ shows a counterfactual scene, which contradicts the world knowledge or common sense (e.g., an image of the Statue of Liberty holding a sword). The reason for using counterfactual images is that the VP needs to evaluate visual perception capabilities on the same images that are used for language prior assessments, where the use of counterfactual images is essential.
@@
[…]
@@ 3.3 Commonsense Bias (CB)
It has been observed that LVLMs, including LLMs, exhibit a reluctance to provide responses that contradict common sense or learned world knowledge, even when they are explicitly instructed to respond based on counterfactual contexts [3, 31, 15, 30]. We propose a Commonsense Bias test (CB) to disentangle this bias from language priors, which is the goal of this benchmark. To eliminate the influence of modality in the evaluation of commonsense bias, we provide LVLMs with a counterfactual textual context $T_{\textrm{cf}}$ and a counterfactual image $I_{\textrm{cf}}$ as input. Also, we provide the models with two statements, $s_{\textrm{cf}}$ and $s_{\textrm{fact}}$, which are true and false respectively under the given context. We wrap the context and statement with a prompt template $\textrm{pr}_{\textrm{CB}}$, which instructs the model to explicitly follow the information provided in the context, rather than common sense.
@@
$\begin{split}P_{\textrm{CB}}=\mathbbm{1}(&\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{CB}}(T_{\textrm{cf}},s_{\textrm{cf}}))=\textrm{``True''}\\ \land\,&\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{CB}}(T_{\textrm{cf}},s_{\textrm{fact}}))=\textrm{``False''}\land P_{\textrm{CK}}=1)\end{split}$ (3)
@@ 3.4 Language Prior (LP)
$\begin{split}P_{\textrm{LP}}=\mathbbm{1}(&\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{LP}}(s_{\textrm{cf}}))=\textrm{``True''}\\ \land\,&\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{LP}}(s_{\textrm{fact}}))=\textrm{``False''}\land P_{\textrm{CB}}=1\land\,P_{\textrm{VP}}=1)\end{split}$ (4)
@@ 4 Data Generation
For each selected concept, we employed GPT-4 (gpt-4-0125-preview) [18] to create 50 instance triples, each consisting of a context, a true statement, and a false statement. We provided a detailed instruction with 3-shot prompt as input, using hand-crafted concept-specific examples to reflect the characteristics of each concept. The examples are designed to be easy in terms of reasoning, to minimize the influence of the models’ reasoning ability and focus solely on measuring language priors.
@@
Counterfactual Images
Next, we proceed with the generation of counterfactual image $I_{\textrm{cf}}$ from the filtered textual contexts. Given the significance of LP in our benchmark, we generate multiple images per test for LP, unlike factual images where we generate only one image per test. We take the average performance on these images, enabling a more accurate evaluation. The images are generated using DALL-E 3 [17], where the textual context $T_{\textrm{cf}}$ is provided as input, and 12 images are sampled. To provide diversity of image style, we produce four images each in photorealistic, illustration, and cartoon styles per one textual context. Consequently, for the 421 contexts, a total of 5,052 images are generated.
Table 1: The number of instance triples and images for each concept.
@@
To verify the quality of the generated $I_{\textrm{fact}}$, $s_{\textrm{exist}}$, and $s_{\textrm{nil}}$, we evaluate whether OpenAI GPT-4o (gpt-4o-2024-05-13) [20], which is the most advanced available LVLM, passes the CK and VP. For instances where GPT-4o fails, human verification was conducted. If the failure was due to an error in the data generation process, we addressed the cause of the error by either regenerating the factual image or manually correcting the nouns in statements.
@@
The statistics of the benchmark data generated through the process are presented in Table 1. The difficulty of data generation varies for each concept, resulting in different proportions of samples being filtered out during the human review process. Ultimately, a total of 302 instance triples and 2,576 images, encompassing both counterfactual and factual images, were included in the benchmark. Data samples for each concept can be found in Appendix B.
@@ 5 Experiments
Table 2: Main experimental results on VLind-Bench.
@@
GPT-4o [20] 93.0 96.0 96.8 89.8 97.0 89.4
GPT-4V [19] 90.1 85.4 90.8 77.6 91.1 75.6
Gemini Pro Vision [7] 80.5 90.4 77.0 79.0 75.5 65.5
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) [10] 94.4 95.7 76.1 58.6 75.5 46.7
LLaVA-NEXT 34B (Nous Hermes 2 Yi 34B) [10] 80.5 85.8 61.7 61.1 67.2 44.5
LLaVA-1.5 13B (Vicuna v1.5 13B) [13] 59.9 92.1 40.9 42.0 31.5 20.9
LLaVA-1.5 7B (Vicuna v1.5 7B) [13] 0.0 0.0 - - 0.0 0.0
+ RLAIF-V [29] 17.9 8.3 48.1 25.0 54.3 35.7
InstructBLIP 13B [5] 66.6 79.5 54.2 57.8 46.7 28.0
InstructBLIP 7B [5] 58.6 73.5 28.2 14.6 27.2 21.0
OmniLMM 12B (Zephyr 7B $\beta$) [28] 88.1 97.7 78.6 81.4 79.5 66.4
MiniCPM-V-2 2.8B [28] 76.2 98.3 56.5 68.1 49.0 34.1
@@
Qwen 1.5 72B Chat [2] 75.8 - 69.9 - 74.2 -
Nous Hermes 2 Yi 34B [16] 83.1 - 75.3 - 77.8 -
Vicuna v1.5 13B [22] 57.9 - 80.0 - 69.2 -
Vicuna v1.5 7B [22] 0.0 - - - 0.0 -
Zephyr 7B $\beta$ [23] 62.3 - 45.7 - 40.7 -

### v4

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated outstanding performance across various multimodal tasks. However, they suffer from a problem known as language prior, where responses are generated based solely on textual patterns while disregarding image information. Addressing the issue of language prior is crucial, as it can lead to undesirable biases or hallucinations when dealing with images that are out of training distribution. Despite its importance, current methods for accurately measuring language priors in LVLMs are poorly studied. Although existing benchmarks based on counterfactual or out-of-distribution images can partially be used to measure language priors, they fail to disentangle language priors from other confounding factors. To this end, we propose a new benchmark called VLind-Bench, which is the first benchmark specifically designed to measure the language priors, or blindness, of LVLMs. It not only includes tests on counterfactual images to assess language priors but also involves a series of tests to evaluate more basic capabilities such as commonsense knowledge, visual perception, and commonsense biases. For each instance in our benchmark, we ensure that all these basic tests are passed before evaluating the language priors, thereby minimizing the influence of other factors on the assessment. The evaluation and analysis of recent LVLMs in our benchmark reveal that almost all models exhibit a significant reliance on language priors, presenting a strong challenge in the field. [^1] Evaluation code and benchmark data are available at https://github.com/klee972/vlind-bench.
VLind-Bench: Measuring Language Priors in Large Vision-Language Models
[…]
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across various tasks through pre-training on massive multimodal datasets and visual instruction tuning. (Liu et al., 2023; Dai et al., 2023a; Zhu et al., 2024; Ye et al., 2024; Peng et al., 2023). However, these models tend to generate responses based solely on spurious text patterns, leaving the given image unconsidered. We refer to this problem as language prior, borrowing the term from the Visual Question Answering (VQA) community (Agrawal et al., 2018). Such language priors can lead to undesirable biases (Hall et al., 2023) and hallucinations (Wang et al., 2023). For example, when a model is presented with an image of a red banana and a yellow apple along with the question, “Is the banana yellow?,” it has been observed that the model frequently responds with “Yes,” ignoring the image content (Zhou et al., 2023). To develop a trustworthy LVLM, resolving the language prior issue is crucial; however, it has not been explored much nor has benchmarks that can accurately measure the issues.
@@
One approach to measure language priors is assessing performance on VQA benchmarks consisting of counterfactual images (e.g., Whoops! (Bitton-Guetta et al., 2023) and ROME (Zhou et al., 2023)). If a model bears language priors, it will answer the question based on learned facts or common sense from its parametric knowledge without collaborating information in the given context (i.e., image); easily failing on answering counterfactual VQA tasks. However, it is challenging to distinguish the models’ misbehaviors solely caused by language priors from those caused by other deficiencies in LVLMs. For example, there could be multiple factors affecting performance in counterfactual-contents VQA tasks – not only language priors but also commonsense knowledge, visual perception capabilities, and the model’s reluctance to counterfactual responses. Such confounding factors make it difficult to evaluate methodologies for improving language prior problems and to assess progress in the research field.
@@
With VLind-Bench, we evaluate recent open-source and proprietary LVLMs’ language priors. The results show that all of the models except GPT-4o (OpenAI, 2024) suffer from excessive reliance on language priors, demonstrating the challenging nature of our benchmark and the need for further improvements. Furthermore, our experiment and analysis on existing LVLMs show that the influence of language priors is inversely proportional to the scale of the backbone LLM. We also reveal that Reinforcement Learning from Human Feedback (RLHF) techniques (Yu et al., 2024a, b), which are designed to mitigate hallucinations, can help reduce the reliance on language priors.
@@ 2.1 Large Vision-Language Models
Recently, there has been a lot of effort in extending Large Language Models (LLMs) to include visual inputs, forming a new class of models known as Large Vision-Language Models (LVLMs) (Liu et al., 2023; Dai et al., 2023a; Zhu et al., 2024; Ye et al., 2024; OpenAI, 2023b, 2024; Google, 2024). These LVLMs are gaining attention as a new paradigm in vision-language learning by transferring the exceptional properties of LLMs, such as multi-step reasoning ability and in-context learning, to the multimodal domain. However, these LVLMs are not free from the bias and hallucination issues inherent in LLMs (Hall et al., 2023; Li et al., 2023; Gunjal et al., 2024; Zhou et al., 2024a; Dai et al., 2023b; Min et al., 2024). Despite this, creating benchmarks to diagnose these problems is more challenging with the image modality, leading to slower progress in benchmark development compared to LLMs.
@@ 2.2 Benchmarks with Counterfactual Context
Since counterfactual contexts can assess the robustness and generalization capabilities of LLMs or LVLMs, several benchmarks utilizing this approach have been proposed. These benchmarks assume that if a model responds based on memorized facts without properly understanding the context of text or images, it would fail to correctly solve tasks conditioned on counterfactual contexts. Benchmarks such as IfQA (Yu et al., 2023) and DisentQA (Neeman et al., 2023) counterfactually augment textual contexts to determine whether the language model accurately incorporates augmented information when answering questions. Wu et al. (2024) evaluate LLMs on reasoning tasks based on counterfactual contexts. Benchmarks like Whoops! (Bitton-Guetta et al., 2023), ROME (Zhou et al., 2023), HallusionBench (Guan et al., 2024), and ViLP (Luo et al., 2024) evaluate the counterfactual reasoning abilities of multimodal models by conducting VQA tasks conditioned on counterfactual images. However, these benchmarks cannot disentangle the reliance on language priors and commonsense biases of a model.
@@ 3 Benchmark Structure
If we introspect how a human solves an counterfactual vision-language task, the three additional assessment types we propose appear more convincing. Humans first understand the image through visual perception, then retrieve real-world information about the objects using commonsense knowledge, and finally reason about how the given situation deviates from real-world common sense. Decomposing multimodal counterfactual reasoning into these three steps is a very natural approach, and each of these steps directly corresponds to our Visual Perception, Commonsense Knowledge, and Commonsense Bias tasks.
@@ 3.1 Commonsense Knowledge (CK)
$\begin{split}P_{\textrm{CK}}=&\mathbbm{1}(\textrm{LVLM}(I_{\textrm{fact}},\textrm{pr}_{\textrm{CK}}(s_{\textrm{fact}}))=\textrm{``True''}\\ &\land\textrm{LVLM}(I_{\textrm{fact}},\textrm{pr}_{\textrm{CK}}(s_{\textrm{cf}}))=\textrm{``False''}),\end{split}$ (1)
@@ 3.2 Visual Perception (VP)
The fundamental ability underpinning all multimodal tasks is visual perception, particularly the ability to recognize objects (Locatello et al., 2020; Burgess et al., 2019). Similar to the CK, evaluating a model on more complex tasks would be meaningless when it fails in object recognition. Therefore, we introduce the Visual Perception test (VP) to assess whether LVLMs can recognize objects in a given counterfactual image. VP consists of one counterfactual image $I_{\textrm{cf}}$ and two statements $s_{\textrm{exist}}$ and $s_{\textrm{nil}}$. Contrary to the CK, the image $I_{\textrm{cf}}$ shows a counterfactual scene, which contradicts the world knowledge or common sense (e.g., an image of the Statue of Liberty holding a sword). The reason for using counterfactual images is that the VP needs to evaluate visual perception capabilities on the same images that are used for language prior assessments, where the use of counterfactual images is essential.
@@
$\begin{split}P_{\textrm{VP}}=&\mathbbm{1}(\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{VP}}(s_{\textrm{exist}}))=\textrm{``True''}\\ &\land\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{VP}}(s_{\textrm{nil}}))=\textrm{``False''})\end{split}$ (2)
@@ 3.3 Commonsense Bias (CB)
It has been observed that LVLMs, including LLMs, exhibit a reluctance to provide responses that contradict common sense or learned world knowledge, even when they are explicitly instructed to respond based on counterfactual contexts (Bitton-Guetta et al., 2023; Zhou et al., 2023; Neeman et al., 2023; Yu et al., 2023). We propose a Commonsense Bias test (CB) to disentangle this bias from language priors, which is the goal of this benchmark. To eliminate the influence of modality in the evaluation of commonsense bias, we provide LVLMs with a counterfactual textual context $T_{\textrm{cf}}$ and a counterfactual image $I_{\textrm{cf}}$ as input. Also, we provide the models with two statements, $s_{\textrm{cf}}$ and $s_{\textrm{fact}}$, which are true and false respectively under the given context. We wrap the context and statement with a prompt template $\textrm{pr}_{\textrm{CB}}$, which instructs the model to explicitly follow the information provided in the context, rather than common sense.
@@
$\begin{split}P_{\textrm{CB}}=&\mathbbm{1}(\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{CB}}(T_{\textrm{cf}},s_{\textrm{cf}}))=\textrm{``True''}\\ &\land\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{CB}}(T_{\textrm{cf}},s_{\textrm{fact}}))=\textrm{``False''}\\ &\land P_{\textrm{CK}}=1)\end{split}$ (3)
@@ 3.4 Language Prior (LP)
$\begin{split}P_{\textrm{LP}}=&\mathbbm{1}(\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{LP}}(s_{\textrm{cf}}))=\textrm{``True''}\\ &\land\textrm{LVLM}(I_{\textrm{cf}},\textrm{pr}_{\textrm{LP}}(s_{\textrm{fact}}))=\textrm{``False''}\\ &\land P_{\textrm{CB}}=1\land\ P_{\textrm{VP}}=1)\end{split}$ (4)
@@ 4 Data Generation
For each selected concept, we employed GPT-4 (gpt-4-0125-preview) (OpenAI, 2023a) to create 50 instance triples, each consisting of a context, a true statement, and a false statement. We provided a detailed instruction with 3-shot prompt as input, using hand-crafted concept-specific examples to reflect the characteristics of each concept.
@@
Table 1: The number of instance triples and images for each concept.
Dataset Num. category/tags Num. images Num. image-question pairs
WHOOPS! Bitton-Guetta et al. (2023) 26 500 10,874
ROME Zhou et al. (2023) 5 1,563 10,941
IfQA Yu et al. (2023) 7 - 6,606
VLind-Bench 11 2,576 14,248
Table 2: Dataset size comparison with similar counterfactual benchmarks.
Counterfactual Images
Next, we proceed with the generation of counterfactual image $I_{\textrm{cf}}$ from the filtered textual contexts. Given the significance of LP in our benchmark, we generate multiple images per test for LP, unlike factual images where we generate only one image per test. The images are generated using DALL-E 3 (OpenAI, 2023c), where the textual context $T_{\textrm{cf}}$ is provided as input, and 12 images are sampled. To provide diversity of image style, we produce four images each in photorealistic, illustration, and cartoon styles per one textual context. Consequently, for the 421 contexts, a total of 5,052 images are generated.
@@
To verify the quality of the generated $I_{\textrm{fact}}$, $s_{\textrm{exist}}$, and $s_{\textrm{nil}}$, we evaluate whether OpenAI GPT-4o (gpt-4o-2024-05-13) (OpenAI, 2024), which is the most advanced available LVLM, passes the CK and VP. For instances where GPT-4o fails, human verification was conducted. If the failure was due to an error in the data generation process, we addressed the cause of the error by either regenerating the factual image or manually correcting the nouns in statements.
@@
The statistics of the benchmark data generated through the process are presented in Table 1. The difficulty of data generation varies for each concept, resulting in different proportions of samples being filtered out during the human review process. Ultimately, a total of 302 instance triples and 2,576 images, encompassing both counterfactual and factual images, were included in the benchmark. We compare the size of VLind-Bench with other counterfactual benchmarks in Table 2. Data samples for each concept can be found in Appendix E.
@@ 5 Experiments
GPT-4o (OpenAI, 2024) 93.0 96.0 96.8 89.8 97.0 89.4
GPT-4V (OpenAI, 2023b) 90.1 85.4 90.8 77.6 91.1 75.6
Gemini Pro Vision (Google, 2024) 80.5 90.4 77.0 79.0 75.5 65.5
@@
LLaVA-NEXT 72B (Qwen 1.5 72B Chat) (Li et al., 2024) 94.4 95.7 76.1 58.6 75.5 46.7
LLaVA-NEXT 34B (Nous Hermes 2 Yi 34B) (Li et al., 2024) 80.5 85.8 61.7 61.1 67.2 44.5
LLaVA-1.5 13B (Vicuna v1.5 13B) (Liu et al., 2024) 59.9 92.1 40.9 42.0 31.5 20.9
LLaVA-1.5 7B (Vicuna v1.5 7B) (Liu et al., 2024) 0.0 0.0 - - 0.0 0.0
+ RLAIF-V (Yu et al., 2024b) 17.9 8.3 48.1 25.0 54.3 35.7
InstructBLIP 13B (Dai et al., 2023a) 66.6 79.5 54.2 57.8 46.7 28.0
InstructBLIP 7B (Dai et al., 2023a) 58.6 73.5 28.2 14.6 27.2 21.0
OmniLMM 12B (Zephyr 7B $\beta$) (Yu et al., 2024a) 88.1 97.7 78.6 81.4 79.5 66.4
MiniCPM-V-2 2.8B (Yu et al., 2024a) 76.2 98.3 56.5 68.1 49.0 34.1
@@
Qwen 1.5 72B Chat (Bai et al., 2023) 75.8 - 69.9 - 74.2 -
Nous Hermes 2 Yi 34B (NousResearch, 2023) 83.1 - 75.3 - 77.8 -
Vicuna v1.5 13B (Team, 2023) 57.9 - 80.0 - 69.2 -
Vicuna v1.5 7B (Team, 2023) 0.0 - - - 0.0 -
Zephyr 7B $\beta$ (Tunstall et al., 2023) 62.3 - 45.7 - 40.7 -
Table 3: Main experimental results on VLind-Bench.